OpenStudy (anonymous):

The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 5y – z = 6 3x – y – 2z = 4 3x – 2y + 3z = 13 a. x + 5y + 6 b. x + 5y – 6 c. –x – 5y + 6 d. –x – 5y – 6

OpenStudy (anonymous):

The question is asking you to solve for z in eqn 1. Can you do that?

OpenStudy (anonymous):

i dont know how to

OpenStudy (anonymous):

x + 5y – z = 6 that is the first equation. Solving for z means that we want z to be by itself. One thing we can do is to add z to both sides. Can you do that?

OpenStudy (anonymous):

x + 5y = z + 6

OpenStudy (anonymous):

Great now we want to subtract 6 from both sides x + 5y = z + 6 x + 5y-6 = z + 6-6 x + 5y-6 = z Can you tell me which answer has this expression x + 5y-6

OpenStudy (anonymous):

answer choice B

OpenStudy (anonymous):

good job